OpenStudy (anonymous):

(integral from 1 to infinity) [ln(x)]/(x)

OpenStudy (freckles):

try the sub u=ln(x)

OpenStudy (anonymous):

I want to know if it converges or diverges

OpenStudy (freckles):

Well we need to integrate first

OpenStudy (anonymous):

yeah the antiderivative i got was (1/2)(ln(x))^2 and then I solved it as the limit as b approaches infinity evaluating it from 1 to b. The answer i got was that it diverges but I'm not sure if that is right.

OpenStudy (freckles):

\[\lim_{b \rightarrow \infty}(\frac{\ln(b))^2}{2}-\frac{\ln(1))^2}{2})=\lim_{b \rightarrow \infty}\frac{(\ln(b))^2}{2}\]

OpenStudy (freckles):

well ln(b)->inf as b->inf

OpenStudy (freckles):

squaring that gets you bigger numbers quicker

OpenStudy (freckles):

I agree with the diverges

OpenStudy (anonymous):

yeeeeesssssssssss!!! thank you so much!